AMS Author Handbook

Contents

General Instructions for Preparing Manuscripts 1 Introduction 1 The four types of submissions 1 Reimbursement 2 Permissions 2 General resources 3 Graphics 3 Equations 4 Submission 4

General Checklist for Authors 5

Information Specific to Monographs 6

Information Specific to Proceedings Volumes 7

Information Specific to Journals 8

Camera-Ready Submissions 9 Camera-ready electronic submissions 9 Camera-ready nonelectronic submissions 10

TEX Resources 12

Author Package Requirements 13

Obtaining and Processing Author Packages from e-MATH 14 How to log onto e-MATH 14 Which files you need and how to get them 15 How to end the FTP session and log off 15 Electronic file installation and processing overview 15

How to Submit Electronic Files and Graphics 16 Electronic mail submissions 16 FTP submissions 16 Diskette submissions 17 Sample FTP and e-mail sessions 17

Checklist for Electronic Submissions 19 Camera-ready and electronic manuscripts 19 Electronic manuscripts only 19 Camera-ready submissions only 19

Quick Start for Obtaining and Using Author Packages 21

List of AMS-TEX commands 22

List of AMS-L^{A}TEX commands 23

References 24

AMS Resources (inside back cover)

Author Packages (back cover)

General Instructions for Preparing Manuscripts

Introduction

This handbook is intended for authors and keyboarders who are preparing books or articles for submission to the American Mathematical Society (AMS).

The AMS, one of the world’s largest publishers of research literature in the mathematical sciences, uses TEX to typeset its publications. You will find general information about TEX in the TEX Resources chapter. Information for contacting the appropriate departments at the AMS appears on the inside back cover of this handbook.

The AMS offers TEX-related support software, such as AMS-TEX and AMS-L^{A}TEX which provide easy ways to code manuscripts that contain mathematical expressions. To encourage the use of AMS-TEX and AMS-L^{A}TEX, the AMS has developed publication-specific author packages. These packages contain instructions and style files. By using the style files, you will shorten the time needed to publish your work and help reduce the price of the publication. You will find instructions for obtaining author packages later in this handbook. They are available free of charge via FTP.

There is a Quick Start chapter which offers step-by-step instructions for obtaining and using author packages.

Additional instructions for correctly encoding AMS-TEX or AMS-L^{A}TEX submissions are given in the Instructions for Preparation of Papers and Monographs (this will be referred to throughout this handbook as the Instructions). These instructions are available electronically in the author package directories on the AMS internet node e-MATH (see Obtaining and Processing Author Packages from e-MATH ) or in printed form from our Customer Services Department.

The four types of submissions

The AMS divides submissions into four types depending upon whether the submission is camera-ready or not and upon whether the submission is electronic or not. In this handbook you will find a general overview of electronic submissions, general information about the different types of TEX and the author packages, a description of resources available, and complete information for the preparation of nonelectronic submissions.

• An electronic manuscript is a TEX file prepared using AMS-TEX or AMS-L^{A}TEX together with the appropriate style file, and following AMS instructions. The file is processed through the AMS production system, and an editorial proofreading is done. In most cases you will receive galley pages to proofread before the final copy is printed.

• A camera-ready electronic submission is a DVI or suitable PostScript file you produce. (Contact AMS Technical Support for further details if you want to submit your entire article in PostScript format.) The final copy is printed by the AMS directly from your file. See the instructions for preparing camera-ready submissions later in this handbook and the Instructions document included in the author package. You are responsible for final formatting and proofreading.

• A nonelectronic manuscript is a manuscript prepared without using the style files. A nonelectronic manuscript is keyed by the AMS as a TEX file. An editorial proofreading is done, and usually a proof is sent to authors before the final copy is printed. For nonelectronic manuscripts it is important that the manuscript be typed with at least double spacing and with generous margins on both sides (1\frac{1}{4} inches). The double spacing provides clear separation of subscripts and superscripts and also permits changes to be inserted directly above portions crossed out. The margins allow the copyeditor to write instructions to the compositor. The pages of the manuscript should be numbered consecutively in the upper right-hand corners. This ensures that the absence of a page or a misplaced page will be noticed.

• A camera-ready nonelectronic submission is material submitted by authors that is photographed by the AMS and then published. See the instructions for preparing camera-ready submissions later in this handbook. You are responsible for final formatting and proofreading.

The types of submissions that can be accepted depend on the specific publication; for example, journals generally do not accept camera-ready submissions. If you do not know the types of submissions accepted by the publication that you are submitting to, contact the Publication Division.

The AMS has created style files which can be used if you are preparing a paper or monograph and are uncertain to which publication you will be submitting. See the list of available author packages on the back cover of this handbook.

Reimbursement

The AMS will, upon request, reimburse authors for actual expenses incurred in the preparation of a camera-ready submission of a monograph designated by the AMS for camera-ready submission or an article in a proceedings volume. All requests must be received with the camera copy. The maximum rate of reimbursement is $8.00 per page. A fee for any work required at the AMS will be deducted from the total amount of the reimbursement. You will find an Author Reimbursement form and a W-9 form included in the Packet for Camera-Ready Submissions (available from the Publication Division). Include these forms with your submission of camera copy. Reimbursement payments are mailed after the book is published.

Permissions

Authors are responsible for obtaining written permission to use another person’s work—including pictures, figures, and other art—whether or not it has been published previously. If the work has been published previously, the publisher is usually empowered to grant permission, but that is not always the case; the author or an heir sometimes retains such rights. Permission is not necessary, of course, for use of work in the public domain (such as U.S. government publications or works on which the copyright has expired), but credit should be given, in any case, to the originator of the work.

Authors are advised to begin writing for permissions as soon as a manuscript has been accepted for publication by an editor or editorial committee. The Publication Division will assist an author in obtaining permissions if such assistance is requested at least six weeks before the paper is due at the AMS. A copy of the letter(s) granting permission must be sent to the AMS with the submission.

General resources

For general advice on writing mathematics, see [SHSD] and [NJH]. For information of a more specific nature including selecting notation, rules for spacing and breaking mathematical expressions, English usage, the appropriate format for references, a list of mathematical symbols, and the symbols used in proofreading, see [MIT] and [NJH]. In your bibliography, use the abbreviations given in [ASMR].

Graphics

Graphics are classified as either electronic or nonelectronic. Refer to the Graphics section of the Instructions document included in your author package if you are uncertain about specific coding. If your submission is prepared as a camera-ready submission, refer to the section on Graphics placement in the Camera-Ready Submissions chapter in this handbook.

Electronic graphics. Figures may be submitted to the AMS in an electronic format. The AMS recommends that graphics created electronically be saved in encapsulated PostScript (EPS) format. This includes graphics originated via a graphics application as well as scanned photographs or other computer-generated images.

Many popular graphics applications under a Macintosh, Windows, or Unix environment allow files to be saved in EPS format. However, if your package does not support EPS output, save your graphics file in one of the standard graphics formats—such as TIFF, PICT, GIF, etc.—rather than in an application-dependent format. For example, if you are using SuperPaint on a Macintosh, do not send files in SuperPaint format. Instead, save the file in PICT format from SuperPaint and send the PICT files to the AMS. Graphics files submitted in an applicationdependent format are not likely to be used. No matter what method was used to produce the graphic, it is necessary to provide a paper copy to the AMS.

Note: Authors using graphics packages for the creation of electronic art should also avoid the use of any lines thinner than 0.5 points in width. Many graphics packages allow the user to specify a “hairline” for a very thin line. Hairlines often look acceptable when proofed on a typical laser printer. However, when produced on a high-resolution laser imagesetter, hairlines become nearly invisible and will be lost entirely in the final printing process.

Screens should be set to values between 15% and 85%. Screens which fall outside of this range are too light or too dark to print correctly.

Nonelectronic graphics. These graphics should be drawn in black ink with clean, unbroken lines on nonabsorbent paper. Whenever possible, fonts used in graphics should match those used in the text. Authors’ original graphics are used whenever possible in AMS publications.

Send the originals of photographs or computer-generated images to the AMS. A photocopy of such an image can be used to identify it. To avoid damage to photographic images, do not use paper clips or staples, and do not tape them to a sheet of paper.

For a color image (whether it is to be converted to a black and white image or is to appear in color), submit one of the following (listed in order of preference):

• print,

• slide,

• transparency.

Equations

Check displayed equations carefully, making sure they are broken and aligned following the guidelines in [MIT, pp. 38–41].

Submission

Electronic submissions. See the chapter How to Submit Electronic Files and Graphics.

Nonelectronic submissions. Send two proof copies of your article or monograph to the appropriate journal or book series editor. Include any artwork.

After acceptance. All articles and monographs must have been accepted for publication by the appropriate editorial board before they are sent to the AMS. Camera-ready submissions should be sent to the Publication Division.

Submissions for book monograph or proceedings volumes should be sent to the attention of either Donna Harmon or Christine Thivierge. All journal submissions should be sent to the attention of Christine Vendettuoli.

General Checklist for Authors

All authors should refer to the following list. Information specific to the different types of publications (i.e., books and journals) is listed on the following three pages.

• Authors should proofread their work before submitting it to the AMS.

• The first page of the article must contain a descriptive title. The descriptive title should be short, but informative; useless or vague phrases such as “some remarks about” or “concerning” should be avoided.

• Give the 1991 Mathematics Subject Classification numbers representing the primary and secondary subjects of the work. A list of these numbers may be found in the annual index of Mathematical Reviews, published with the December issue starting in 1990, as well as from the electronic service

e-MATH (telnet e-math.ams.org or 130.44.1.100; login and password

are e-math) or in printed form from Customer Services.

• Give information on grants or contracts under which the research was performed, including grant number.

• An abstract may comprise multiple paragraphs and include displayed material if appropriate. The length of the abstract depends primarily on the length of the paper itself and on the difficulty of summarizing the material, but an upper limit of about 150 words for short papers and 300 words for long papers is suggested. An abstract is not required for book proceedings.

Information Specific to Monographs

• In chapter titles, the first and last words of the title and all nouns, pronouns, adjectives, adverbs, and verbs should be capitalized; articles, conjunctions, and prepositions should be lowercased except for the first and last words of the title.

• The running heads on the left-hand (even-numbered) pages should have the chapter title in uppercase letters. The title in running heads should be shortened only if necessary to fit on one line, leaving room for the page number. The running heads on the right-hand (odd-numbered) pages should have the main-heading title (shortened if necessary) in uppercase letters.

• The copyright page should include primary classification using the the 1991 Mathematics Subject Classification scheme, grant information, and an abstract. For translations the copyright page should include the name and city of the original publisher, the copyright year of the original publication, and the name and address of the translator.

Information Specific to Proceedings Volumes

• In article titles, the first and last words of the title and all nouns, pronouns, adjectives, adverbs, and verbs should be capitalized; articles, conjunctions, and prepositions should be lowercased except for the first and last words of the title.

• The running heads on the left-hand (even-numbered) pages should have the name(s) of the author(s) in uppercase letters. Names in running heads should match the names as given on the first page and should be shortened only if necessary to fit on one line, leaving room for the page number. The running heads on the right-hand (odd-numbered) pages should have the article title (shortened if necessary) in uppercase letters.

• Preliminary papers, as well as abstracts and summaries, are not accepted for publication in AMS proceedings volumes.

• Unmarked, unnumbered footnotes on the first page of an article should include primary classification using the 1991 Mathematics Subject Classification scheme followed by grant information.

• Academic or other affiliations should appear at the end of your article, after the bibliography or references. A current address, if different from the affiliation, should follow the affiliation on a new line with a blank line in between. An e-mail address should be included if available, on a new line following the current address with a blank line in between. Note: Line spacing is taken care of automatically by AMS style files.

• A Copyright Transfer Agreement is sent to the author(s) when the final paper is received at the AMS. Authors retain the right to use all or part of their own work in future publications of their own. They are, however, asked, but not required, to sign other rights over to the AMS. The form should be returned to the Production Department immediately, either signed or with a note that the author(s) will hold copyright.

Information Specific to Journals

• In article titles only the first word and proper nouns should be capitalized.

• The running heads on the left-hand (even-numbered) pages should have the name(s) of the author(s) in uppercase letters. Names in running heads should match the names as given on the first page and should be shortened only if necessary to fit on one line, leaving room for the page number. The running heads on the right-hand (odd-numbered) pages should have the article title (shortened if necessary) in uppercase letters.

• Transactions, Proceedings, Mathematics of Computation, and Bulletin articles (except Book Reviews) must contain an abstract. The main purpose of the abstract is to enable readers to take in the nature and results of the article quickly. Zentralblatt publishes authors’ abstracts instead of reviews, so the abstract may also appear there.

• Unmarked, unnumbered footnotes on the first page of an article should include primary classification according to the 1991 Mathematics Subject Classification scheme followed by grant information. A footnote of key words and phrases, describing the subject matter of the article and taken from it, is optional.

• Academic or other affiliations should appear at the end of your work, after the bibliography or references. A current address, if different from the affiliation, should follow the affiliation on a new line with a blank line in between. An e-mail address should be included if available. Note: Line spacing is taken care of automatically by AMS style files.

• A Copyright Transfer Agreement is sent to the author(s) when the final paper is received at the AMS. Authors retain the right to use all or part of their own work in future publications of their own. They are, however, asked, but not required, to sign other rights over to the AMS. The form should be returned to the Production Department immediately, either signed or with a note that the author(s) will hold copyright.

Camera-Ready Submissions

You should obtain a Packet for Camera-Ready Submissions from the Publication Division. The packet contains a copy of this handbook, model paper (for authors preparing a nonelectronic submission), and reimbursement forms. You may send a few sample pages to the Production Department to check before finalizing copy. Since camera-ready submissions are printed directly from the material supplied by authors, you are responsible for final formatting and proofreading.

Camera-ready electronic submissions

To encourage the use of AMS-TEX or AMS-L^{A}TEX, the AMS has developed publication-specific author packages. These packages contain instructions and style files. By using the style files, you will shorten the time needed to publish your work and help reduce the price of the publication. You will find instructions for obtaining author packages later in this handbook. Author packages are available free of charge via FTP.

Guidelines for paging. Each AMS style file contains macros which set most of the paging standards. Page dimensions, heading formats, and penalties which prevent widows and orphans are just some of the specifications you will not need to worry about when using the style file. Refer to the Instructions for additional information.

Graphics placement. Graphics can be embedded in the output or white space can be left for graphics to be stripped in later. In either case, they should

• be numbered consistently throughout the paper,

• be placed at the top or bottom of the page, and

• have an in-text reference.

You should always use the predefined figure macros to include graphics (refer to the Instructions for specific coding instructions). These macros allow the user to either include encapsulated PostScript (EPS) graphics or specify white space for graphics to be stripped in later. Extra space surrounding the caption and the actual figure is built into the macro. Therefore, if white space is placed where a figure will be set, only the actual size of the figure should be given within the macro.

Make sure all graphics do not extend into the margins.

Ideally, a figure or table should not precede its first text reference unless the graphic and its text reference appear on facing pages (an even-numbered page followed by odd-numbered page), and a figure or table must definitely appear within the same section as its first text reference. When a figure or table is an integral part of text, it may appear unnumbered in place in the middle of text.

Embedded graphics. There are two ways of utilizing EPS graphics with a TEX document:

1. Calls to EPS files can be embedded within the TEX file and automatically incorporated by TEX.

2. Blank space of an appropriate size for each graphic can be left in the TEX document. Graphics files can then be sent to the AMS for high-resolution typesetting and manual positioning in the typeset document.

In either case, submit a separate file for each graphic along with the TEX document. In cases where files cannot be saved in an EPS format, you may choose to leave space in the TEX document and submit the graphics files separately.

If you choose to submit a file with embedded references to external EPS files, incorporate the EPS files into the TEX file using one of the following public-domain macro packages: epsf.tex (or epsf.sty), developed by Radical Eye Software;

psfig.tex, version 1.9 or later; or boxedeps.tex (or boxedeps.sty). The AMS

does not provide these macro files to authors, as they are widely available in the TEX community. For a list of public domain archives contact the Electronic Products and Services Department at the AMS. Also, there are a few requirements if these macros are used:

1. DO NOT include any path names of the included PostScript files. For in-

      stance, do not say \psfigfile=/usr/joe/book/figures/fig1.ps. Instead

remove all explicit path references, so that the above example would become

      \psfigfile=fig1.ps.

2. Include printed copies of all of the PostScript files for the graphics with the TEX file.

3. If you are preparing a camera-ready submission and you are using psfig.tex, use the version which is compatible with dvips, rather than the OzTEX version of that file. If you use boxedeps.tex, when you run TEX to create the DVI or PostScript file to send to the AMS, you must place the command

      \SetRokickiEPSFSpecial immediately after the file boxedeps.tex input.

Camera-ready nonelectronic submissions

A camera-ready nonelectronic submission need not be printed on model paper, but the instructions that appear on the model paper must be followed in order to maintain uniformity within a volume and within a book series. It is particularly important that the type block measure 30 by 50.5 picas (approximately 5 × 8\frac{3}{8} inches). For proceedings articles, the first page, including title, should measure only 36\frac{1}{2} picas (approximately 6 inches) in length. Similarly, in a monograph, the opening page of each chapter should measure 30 by 43\frac{1}{2} picas (approximately 5×7\frac{1}{4} inches). For all other pages, running heads should be included in the measurement of the type block and should be centered at the top of each page in small capital letters.

Guidelines for paging. Avoid the following situations whenever possible:

• A widow, which is a short line at the top of the page that ends a paragraph and is less than half the text width. Use your discretion for lines that approach half the text width. Transfer these lines to the previous page, or transfer a line from the previous page forward.

• A centered heading at the bottom of a page with less than 2 lines of text following it or at the top of a page with only 1 line of text preceding it. (A centered heading can appear at the top of a page. For headings that run into text, there should be at least 2 lines of text, following the heading, at the bottom of the page.)

• The first line of a theorem, definition, remark, etc., at the bottom of a page. If the last line of a theorem, definition, remark, etc., is a widow, follow the rules for widows.

• An unacceptable math break in text between pages. Page breaks for math follow the same guidelines as line breaks for math; see [MIT, pp. 36–37].

• A hyphenated word at the bottom of a right-hand page.

For proceedings articles, aim for an acceptable-looking last page for each article. Avoid having just one reference, the affiliation, current address, e-mail address, or “translated by” line standing alone on the page.

Graphics placement. Graphics can be embedded in the output or white space can be left for graphics to be stripped in later. In either case, they should

• be numbered consistently throughout the paper,

• be placed at the top or bottom of the page, and

• have an in-text reference.

If you are leaving white space for graphics that will be stripped in and you are not using an AMS style file, then you must leave space for graphics as follows:

• at least 1.5 picas (\frac{1}{4} inch) of space between the text and the top of the graphic (if the graphic follows text),

• exact space for the graphic,

• exactly 1 pica of space from the bottom of the graphic to its caption,

• at least 1.5 picas (\frac{1}{4} inch) of space from the caption to the text below (if text follows the graphic).

A figure or table should not precede its first text reference unless the graphic and its text reference appear on facing pages (an even-numbered page followed by odd-numbered page), and a figure or table must definitely appear within the same section as its first text reference. When a figure or table is an integral part of text, it may appear unnumbered in place in the middle of text.

Footnotes. A footnote must appear on the same page as its text reference. On rare occasions it will happen that a footnote will not fit in its entirety on the same page as its text reference. In this case break the footnote between pages at a line of the footnote that ends in the middle of a sentence. This will signify to the reader that the footnote continues on the next page. Please note that the footnote must at least begin on the same page as its text reference.

TEX Resources

L^{A}TEX and plain TEX are available on the Internet free of charge. There are also numerous commercial TEX packages. Making TEX Work [NM] explains and reviews most of the popular public domain and commercial packages and includes instructions for obtaining these packages.

L^{A}TEX is the most popular of the public domain packages designed for use with plain TEX, the basic TEX typesetting program. Whereas plain TEX defines most of the base macros you will need, L^{A}TEX defines prepackaged stylistic macros. It sets up a book, article, report, and letter style, which you can then alter to your specifications. L^{A}TEX User’s Guide & Reference Manual [LM] is the standard instruction manual for L^{A}TEX. The standard reference for plain TEX is The TEXbook [TB]. Both can be obtained from the AMS Customer Services Department.

The TEX Users Group is your best resource for information on both plain TEX and L^{A}TEX. They can be contacted at:

TEX Users Group PO Box 869 Santa Barbara, CA 93102 (805)963-1338

   tug@tug.org

AMS-TEX and AMS-L^{A}TEX are macro packages that provide a variety of mathematical environments for the user. The macro package AMS-TEX provides advanced mathematical typesetting capabilities to the plain TEX user. The AMS-L^{A}TEX macro package provides advanced mathematical typesetting capabilities of AMS-TEX to the L^{A}TEX user. Nearly all of the AMS-TEX features such as multiline equations, matrices, math accents, etc., are incorporated into AMS-L^{A}TEX. Included in these packages are style files which reflect the appearance of the AMS’s standard book and journal design.

Both AMS-TEX and AMS-L^{A}TEX require the installation of TEX. AMS-L^{A}TEX requires the additional installation of L^{A}TEX. (See Author Package Requirements section.)

AMS-TEX and AMS-L^{A}TEX are available via anonymous FTP from e-MATH or for a minimal fee from Customer Services Department. (See the chapter Obtaining and Processing Author Packages from e-MATH.) AMS distributions of AMS-TEX 2.1, AMS-L^{A}TEX 1.2, and AMSFonts 2.2 each include a copy of the related User’s Guide. Each guide provides important information including installation procedures and coding examples. An electronic version of each of these guides is available on e-MATH. Printed versions of each User’s Guide may be ordered free of charge from Customer Services.

Author Package Requirements

The author packages developed by the AMS produce the style of a specific publication using a standard set of command sequences (see the lists of commands later in this handbook). The AMS strongly encourages authors to prepare their articles and monographs using the appropriate author package. By using the appropriate author package, you will shorten the time needed to publish your work and help reduce the price of the publication.

Author packages contain sample paper(s), templates for coding topmatter, an electronic copy of this handbook, instructions for using the package, and style files that produce the style of a specific publication. Instructions for obtaining author packages are included later in this handbook.

Each author package is designed to work with the most recent versions of the base software. If you use older versions of TEX, L^{A}TEX, or AMS-created macro packages, you will run into problems such as bugs which have been fixed in a later version of the software you are running.

To use an author package, you will need the current version of either AMS-TEX or AMS-L^{A}TEX.

If you are preparing an electronic submission using AMS-TEX, you need:

• plain TEX and Computer Modern fonts,

• AMS-TEX (including amsppt.sty),

• AMSFonts, and

• the appropriate publication-specific AMS-TEX author package.

If you are preparing an electronic submission using AMS-L^{A}TEX, you need:

• TEX and Computer Modern fonts,

• L^{A}TEX2ϵ,

    • AMS-L^{A}TEX (including amsart.cls, amsproc.cls, and amsbook.cls),

• AMSFonts, and

• the appropriate publication-specific AMS-L^{A}TEX author package.

Note: You must have L^{A}TEX2ϵ on your system in order to run AMS-L^{A}TEX 1.2. Current AMS-L^{A}TEX author packages have been created to run only under version 1.2 or later since the release of L^{A}TEX2ϵ.

Obtaining and Processing Author Packages from e-MATH

The most recent list of available author packages is on e-MATH. In this section we give a general introduction to the FTP commands to be used for obtaining author packages. This is followed by specific instructions for obtaining author packages from the AMS’s internet node e-MATH (e-math.ams.org) using a standard file transfer method known as “FTP”.

e-MATH is a Unix environment; therefore file and directory names are casesensitive; directory names are lowercase. The basic FTP commands you will need to know are listed below. Note: Some implementations of FTP function slightly differently. For example, you may have to use quit instead of exit. If you have any problems, contact the support people at your site. If you do not have support help available, send mail to support@e-math.ams.org.

FTP Commands

cd (directory name) connect to a directory on the remote host cdup connect to the parent of current directory on

the remote host

lcd (directory name) locally connect to a directory on your com-

puter

pwd ask for name of current directory on the re-

mote host

lpwd* ask for name of current directory on your com-

puter

dir list contents of current directory on the re-

mote host

ldir* list contents of current directory on your com-

puter

get (file name) get a single file from the remote host mget (file specifications) get multiple files from the remote host binary when you are about to transfer binary files ascii when you are about to transfer text files (most

of the files on e-MATH are text files)

Ctrl a shows the amount transferred Ctrl c cancels a transfer in progress exit leave FTP and return to your system

* These commands may not be valid with all versions of FTP.

How to log onto e-MATH

From your host system account type ftp e-math.ams.org <return>. When you see a message indicating that a connection has been opened, log in using the username anonymous. If your version of FTP is now prompting you for a Name or

Username, simply type anonymous <return>. Otherwise, type login anonymous

<return>. You will be asked to enter a password; enter your e-mail address here and press <return>.

Which files you need and how to get them

On the back cover of this handbook there is a list of publications for which author packages are available. To obtain the most up-to-date list of available packages,

retrieve and read the file /pub/author-info/READ.ME.

If you want to retrieve the author package for the Contemporary Mathematics monograph series for use with AMS-TEX, do the following:

1. Connect to the directory by typing:

      cd /pub/author-info/packages/conm-m/amstex

(Use the codes on the back cover of this handbook to replace conm-m with the code for the specific publication you want; change amstex to amslatex if appropriate.)

2. Enter pwd to make sure that Step 1 was successful.

3. Enter dir to see how many files are in the directory. (optional)

4. Connect to the directory on your computer where the files will go, using the lcd command. (Note: The format in which you type the directory name should be the normal format in which you type it on your operating system.)

5. Enter lpwd to make sure that Step 3 was successful.

6. Enter mget *.*

7. Enter ldir to be sure that all the files were copied. (optional) Note: If you are planning to get both AMS-TEX and AMS-L^{A}TEX versions of

author packages, you must set up separate directories on your system. There are files with similar names between the two types of packages; moving them to the same directory will cause some files to be overwritten.

To obtain the TEX macro packages (AMS-TEX and AMS-L^{A}TEX) and fonts (AMSFonts) from e-MATH, follow the instructions in the READ.ME file in the directory /pub.

Questions concerning the above files can be sent to: tech-support@ math.ams.org. Reports of problems in accessing the e-MATH system itself should

be sent to: support@e-math.ams.org.

How to end the FTP session and log off

To exit FTP, simply type exit or quit <return>.

Electronic file installation and processing overview

The following is a general overview of what to do once you obtain your author package.

• If you use AMS-TEX, install the style files in the directory where you have amsppt.sty. If you use AMS-L^{A}TEX, install the style files in the directory

  where you have amsart.cls, amsproc.cls, and amsbook.cls.)

• Print the READ-*.ME file; TEX and print all *.tex files in the package.

• If you do not know whether your submission should be an electronic manuscript or a camera-ready submission, contact the Publication Division at pub@math.ams.org. Once you verify this information follow the appropriate guidelines in this handbook.

• Proofread your work carefully.

• Submit your files following the instructions in the chapter How to Submit Electronic Files and Graphics.

How to Submit Electronic Files and Graphics

First send two proof copies of the article or monograph to the appropriate journal or book series editor. You should state your intention to submit your work electronically, but the electronic version should not accompany the proof copy. Upon acceptance of the paper, the source file(s) should be sent to the Providence office (this includes the TEX source file, any graphics files, and, if the submission is for a camera-ready series, the DVI or PostScript file).

Before sending the source file, be sure you have proofread your article or monograph carefully. The files that you send must be the EXACT files used to generate the proof copy that was accepted for publication. In order to avoid any possible production problems, before sending the files, be sure to reread the Checklist for Electronic Submissions chapter in this handbook. If your submission is to be camera-ready, reread the section Guidelines for paging and be sure that any necessary forms are sent along with your submission.

For all publications, authors are required to send a printed copy of their paper, which exactly matches the copy approved for publication, along with any graphics that will appear in the paper, to the Publication Division.

TEX files may be submitted by e-mail, FTP, or on diskette. The DVI file(s) (for camera-ready submissions) and PostScript files should be submitted only by FTP or on diskette unless they are encoded properly to submit through e-mail. (DVI files are binary and PostScript files tend to be very large.)

If you created your file(s) using Textures, these files have two components—the TEX source code and the DVI component. You should use DVItool to extract the DVI file. Once you have extracted the DVI file, you can FTP it to the incoming directory on e-MATH. If you are unfamiliar with this process, please send us the file on a Macintosh disk, and we will extract the DVI file for you.

Electronic mail submissions

Files sent by electronic mail should be addressed to the Internet address

pub-submit@math.ams.org.

The subject line of the message should use the publication code to identify the series or journal (see inside back cover for list of codes). For a monograph include the name(s) of the author(s), and for an article in a proceedings include the name(s) of the editor(s). If one file has been broken into parts because of its large size, also identify the part, “part n of m”, in the subject line. By including this information in the subject line, you will help speed up the processing of your submission.

Submissions received through e-mail will be acknowledged upon receipt by an automatic reply while your submission is reviewed. If there are any problems with the file received, you will be notified. Do not send DVI files by electronic mail unless they have been appropriately encoded for this type of transmission.

FTP submissions

TEX source files, DVI files, and PostScript files can be transferred over the Internet by FTP to the Internet node e-math.ams.org (130.44.1.100) using the following instructions.

Type ftp e-math.ams.org <return>. When you see a message indicating that a connection has been opened, log in using the username anonymous. If your version of FTP is now prompting you for a Name or Username, simply type

anonymous <return>. Otherwise, type login anonymous <return>. You will be

asked to enter a password; enter your e-mail address here and press <return>.

Once logged in, connect down to the directory “incoming”. This is the only directory in the FTP area that is open to write access. Be sure to set your mode to binary before you transfer the DVI or PostScript file (see sample FTP session on next page). If a file already exists by the name that you choose, you will not be able to write to this directory. After the transfer is complete, you must send mail to pub-submit@math.ams.org to inform the AMS staff of the transfer. In your message tell us the name of the file(s) you transferred, the publication it is intended for, and the name of the editor of the volume (if applicable).

Diskette submissions

TEX source files, DVI files, and PostScript files can be sent to the Publication Division on diskettes. Diskettes of any standard format and density appropriate for either IBM PC-compatible or Macintosh are acceptable. Diskettes should be labelled with your name, the title of your paper or monograph, and the publication for which it is intended.

If you are reusing a diskette, make sure that all irrelevant files that may remain from other work have been deleted. Time spent by the production staff in trying to identify extra files will delay processing of your paper.

A package containing a diskette should be clearly marked “Floppy disk. Do not fold or bend. Do not expose to electric motors or other magnetic fields.” Mailers designed for diskette shipment are available from many sources; if you do not use a special mailer, cardboard or some other stiffener should be enclosed in the package for protection.

Files received on diskette will be acknowledged as soon as possible after they are received. If you have not received an acknowledgment within a reasonable amount of time (two weeks if you are located in North America; three weeks if in Europe, Australia, Japan, or South America; a month if you are in Eastern Europe or Asia), please inquire, writing to the same address to which you sent the diskette.

Sample FTP and e-mail sessions

Sample FTP session. Below is a sample FTP session where the files grey-ch1.tex and grey-ch1.dvi are transferred to e-MATH. Computer messages and prompts are shown in italic text, and the commands that you will enter are shown in typewriter text. Note: Different implementations of FTP work slightly differently, so the messages and prompts you see will not necessarily be identical to those listed below.

   sol08⟩ ftp e-math.ams.org

Connected to e-math.ams.org.

220 e-math FTP server (ULTRIX Version 4.1 Fri Jul 24 13:34:28 EDT 1992) ready.

   Name: anonymous

331 Guest login ok, send ident as password.

Password:

230 Guest login ok, access restrictions apply.

ftp⟩ cd incoming

250 CWD command successful.

ftp⟩ put grey-ch1.tex

200 PORT command successful. 150 Opening data connection for grey-ch1.tex (130.44.1.8,3654). 226 Transfer complete. local: grey-ch1.tex remote: grey-ch1.tex 717 bytes sent in 0.1 seconds (7 Kbytes/s)

ftp⟩ binary

200 Type set to I.

ftp⟩ put grey-ch1.dvi

200 PORT command successful. 150 Opening data connection for grey-ch1.dvi (130.44.1.8,3671). 226 Transfer complete. local: grey-ch1.dvi remote: grey-ch1.dvi 44544 bytes sent in 0.36 seconds (1.2e+02 Kbytes/s)

ftp⟩ quit

221 Goodbye.

Sample e-mail session. The final step of the electronic submission process is to send e-mail notifying the AMS that the files have been transferred. The following is a sample subject line and message of an e-mail message notifying the AMS of the submission by FTP of an article to appear in a volume of the Proceedings of Symposia in Pure Mathematics.

Subject:  PSPUM/Greene
I have submitted the following files into /incoming on e-math:
    rgreene.tex
    rgreene.dvi
for PSPUM edited by Robert Greene.

Checklist for Electronic Submissions

Camera-ready and electronic manuscripts

Some basic principles are important for effective handling of electronic submissions. Keep these principles in mind when preparing your files.

• AMS style files contain commands to identify all the important features of mathematical papers and monographs; use them, not commands you define yourself.

• Do not redefine any plain TEX, L^{A}TEX, AMS-TEX, or AMS-L^{A}TEX command.

• Put definitions for frequently occurring phrases or mathematical expressions together in the preamble section, before the start of the text of the manuscript. Once a macro is created for an expression, you should use it for every occurrence of that expression.

• Every file should be able to be processed independently; do not use “input” files, but instead include the contents of any such files in the main source file. AMS-L^{A}TEX monographs are an exception to this policy when driver files are used to produce an entire volume.

• Use TEX coding for special fonts (e.g., boldface or italic) only within the text of the manuscript; do not use such coding for any headings or predefined environments, e.g., theorems, etc.

• Citations in the manuscript should be coded using \cite.

• References should include all available information; use the abbreviations of journals and book series in [ASMR].

• The research address or institutional affiliation, current address (if different), and e-mail address of each author should be included.

• Proofread your paper thoroughly and carefully.

Electronic manuscripts only

• Do not use TEX coding to control line and page breaks. Lines and pages will break differently in the published paper from the way they break in the file you submit. If you insert TEX coding for page and line breaks, it will have to be removed for production. That work could offset any time saved by your keyboarding the manuscript, and any change to your TEX file creates a small chance of additional errors being introduced.

• Likewise, avoid explicit horizontal and vertical spacing commands.

Camera-ready submissions only

• Check this handbook and the Instructions document of your author package to be sure that you have followed the appropriate instructions.

• Check your final output for page and line breaks; make sure that all text and equations are within the margins (see Strategy for line and page breaking in [ALG] if you are using AMS-L^{A}TEX). Enter additional line or page breaks following the guidelines in the section Guidelines for paging.

• Submit your DVI and/or PostScript file(s) along with all related TEX and encapsulated PostScript graphics files. (Contact AMS Technical Support for further details if you want to submit your entire article in PostScript format.) The TEX file will be archived. Your work will be printed directly from your output file.

Quick Start for Obtaining and Using Author Packages

The purpose of this ‘Quick Start’ is to help you get an author package and start using it as quickly as possible. To use an author package, you need to have the most recent version of AMS-TEX or AMS-L^{A}TEX already installed; you can get these by FTP from e-MATH. If you will be getting more than one author package, complete Step 4 before getting another package; otherwise, files may be overwritten in Step 2. In Steps 1 through 4, you will FTP an author package from e-MATH and install it on your system. The lists of AMS-TEX and AMS-L^{A}TEX commands in this Handbook help you use the package to prepare your files so the time needed to publish your work is shortened.

STEP 1: Log into e-MATH

      First, type: ftp e-math.ams.org   <return>

When you see a message indicating that a connection has been opened, you need to log in using the username anonymous. If your version of FTP is now prompting you for a Name or Username, simply type anonymous <return>.

   Otherwise, type login anonymous and hit <return>.

You will be asked to enter a password; enter your name here and press

   <return>.

STEP 2: Get Author Package

1. Connect to the directory by typing:

       cd /pub/author-info/packages/conm-m/amstex

(Use the codes on the back cover of this handbook to replace conm-m with the code for the specific publication you want; change amstex to amslatex if appropriate.)

2. Enter pwd to make sure that Step 1 was successful.

3. Enter dir to see how many files are in the directory. (optional)

4. Connect to the directory on your computer where the files will go, using the lcd command. (Note: The format in which you type the directory name should be the normal format in which you type it on your operating system.)

5. Enter lpwd to make sure that Step 3 was successful.

     6. Enter mget *.*

7. Enter ldir to be sure that all the files were copied. (optional)

STEP 3: Exit e-MATH To exit FTP simply type exit <return>. Questions concerning the above files can be sent to: tech-support@

math.ams.org. Reports of problems in accessing the e-MATH system itself

   should be sent to: support@e-math.ams.org.

STEP 4: Install the style files in the directory where you have amsppt.sty. (If you use AMS-L^{A}TEX, install the style files in the directory where you have

   amsart.cls, amsproc.cls, and amsbook.cls.)

STEP 5: TEX and Print 1. Print the READ-*.ME file.

2. TEX and print all *.tex files in the package.

STEP 6: If you do not know whether your submission is an electronic manuscript or a camera-ready submission, contact the Publication Division at pub@ math.ams.org. Once you obtain this information follow the appropriate guidelines in this handbook.

STEP 7: Read the Checklist for Electronic Submissions in this handbook and the Instructions you printed in Step 5; use the *.top files as templates and the *.tex files as examples.

STEP 8: Coding

• AMS style files contain commands to identify all the important features of mathematical papers and monographs; use them, not commands you define yourself. Quick reference lists of these commands are given on the following two pages.

• Do not redefine any plain TEX, L^{A}TEX, AMS-TEX, or AMS-L^{A}TEX command.

• Put definitions for frequently occurring phrases or mathematical expressions together in the preamble section, before the start of the text of the manuscript. Once a macro is created for an expression, you should use it for every occurrence of that expression.

• Use TEX coding for special fonts (e.g., boldface or italic) only within the text of the manuscript; do not use such coding for any headings or predefined environments, e.g., theorems, etc.

STEP 9: Proofread your work carefully; for camera-ready submissions make sure the printing does not run into the margins.

STEP 10: Read the Checklist for Electronic Submissions in this handbook, to be sure your files are ready to be submitted.

STEP 11: Submit your files following the instructions in How to Submit Electronic Files and Graphics in this handbook.

List of AMS-TEX commands

Code  Comments

\specialhead    \endspecialhead Use only if sections are grouped

into parts.

\head           \endhead  First-level head (sections).
\subhead        \endsubhead  Second-level head (subsections).
\subsubhead     \endsubsubhead Third-level head (subsubsections).
\proclaim{ }    \endproclaim  Theorem, Lemma, Corollary,

Proposition, Conjecture, Criterion, Algorithm.

\demo{ }        \enddemo  Proof.

\remark{ }      \endremark  Remark, Note, Notation, Claim,

Summary, Acknowledgment, Case, Conclusion.

\definition{ }  \enddefinition Definition, Condition, Problem.

\example{ }     \endexample  Example.

\xca{ }         \endxca  Exercises in body of text.
\xcb{ }         \endxcb  Exercises at end of chapter

(available only for monographs).

\topinsert  These four lines of code
\vspace #in  create space for a figure,

\botcaption{ }  \endcaption  print caption,

\endinsert  and close the insert.
\cite{ }  Cite a reference.

\roster         \endroster  Use for lists.

\rom{ }  Use to get roman fences, i.e.,

                                 ( ), [ ], { },

and roman punctuation, i.e.,

: (colon) ; (semicolon) , (comma) '' (quotation marks) in the text of proclamations.

List of AMS-L^{A}TEX commands

Code  Comments
\specialsection{}  Use only if sections are grouped

into parts.

\section{}  First-level head (sections).
\subsection{}  Second-level head (subsections).
\subsubsection{}  Third-level head (subsubsections).
\begin{xca}        \end{xca}  Exercises in body of text.
\begin{xcb}        \end{xcb}  Exercises at end of chapter

(available only for monographs).

\begin{figure}  These four lines of code
\vspace{#in}  create space for a figure,
\caption{}  print caption,
\end{figure}  and close the insert
\cite{}  Cite a reference.

\begin{enumerate}  \end{enumerate} Use for numbered lists.
\begin{itemize}    \end{itemize}  Use for bulleted lists.

\textup{}  Use to get roman fences, i.e.,

                                    ( ), [ ], { },

and roman punctuation, i.e., : (colon) ; (semicolon) , (comma) '' (quotation marks) in the text of proclamations.

Theorems and similar structures are treated as environments in L^{A}TEX. Three different theorem styles are provided by AMS-L^{A}TEX: plain, definition, and remark. By referring to these styles and using the \newtheorem command, an author can build a complement of theorem environments appropriate for any paper or monograph. The use of these commands is described in the AMS-L^{A}TEX User’s Guide [ALG]. All \newtheorem specifications should be included in the preamble.

The following list summarizes the types of structures which are appropriate for use with each style.

plain Theorem, Lemma, Corollary, Proposition, Conjecture,

Criterion, Algorithm

definition Definition, Condition, Problem, Example

remark Remark, Note, Notation, Claim, Summary, Acknowledgment, Case, Conclusion

References

[AFG] AMSFonts version 2.2 user’s guide, Amer. Math. Soc., Providence, RI, 1995.

[ALG] AMS-L^{A}TEX version 1.2 user’s guide, Amer. Math. Soc., Providence, RI, 1995.

[ASMR] Abbreviations of names of serials reviewed in Mathematical Reviews, Amer. Math. Soc., Providence, RI, revised annually.

[ATG] AMS-TEX version 2.1 user’s guide, Amer. Math. Soc., Providence, RI, 1992.

[GMS] Michel Goossens, Frank Mittelbach, and Alexander Samarin, The L^{A}TEX companion, Addison-Wesley, Reading, MA, 1994.

[Joy] M. D. Spivak, The joy of TEX, 2nd revised ed., Amer. Math. Soc., Providence, RI, 1990.

[LM] Leslie Lamport, L^{A}TEX: A document preparation system, 2nd revised ed., Addison-Wesley, Reading, MA, 1994.

[MIT] Ellen E. Swanson, Mathematics into type, Amer. Math. Soc., Providence, RI, 1979.

[NJH] Nicholas J. Higham, Handbook of writing for the mathematical sciences, SIAM, Philadelphia, PA, 1993.

[NM] Norman Walsh, Making TEX Work, O’Reilly & Associates, Inc., Sebastopol, CA, 1994.

[SHSD] Norman E. Steenrod, Paul R. Halmos, Menahem M. Schiffer, and Jean A. Dieudonné, How to write mathematics, 4th printing 1993, Amer. Math. Soc., Providence, RI, 1973.

[TB] Donald E. Knuth, The TEXbook, Addison-Wesley, Reading, MA, 1984.

AMS Resources

All written correspondence should be sent to the appropriate department at the American Mathematical Society:

P. O. Box 6248 or 201 Charles Street
Providence, RI 02940-6248  Providence, RI 02904

or by FAX to 401-331-3842.

Questions concerning difficulties in preparing or submitting an AMS-TEX or AMS-L^{A}TEX electronic submission after it has been accepted for publication by the appropriate editorial board should be directed to:

Technical Support Electronic Products and Services Phone: 800-321-4267 or 401-455-4080

Internet: tech-support@math.ams.org

Questions concerning the type of publication you need to prepare (cameraready or typeset series) or requests for obtaining the Packet for Camera-Ready Submissions should be directed to:

Publication Division Phone: 800-321-4267, ext. 4051 or 401-455-4051

Internet: pub@math.ams.org

Copyright Transfer Agreements should be returned to:

Production Department Internet: production@math.ams.org or FAX to 401-331-3842

The AMS has a variety of products to assist authors preparing electronic submissions including AMS-TEX, AMS-L^{A}TEX, and AMSFonts as well as several books and manuals. If you are interested in receiving a free Electronic Products brochure, or for information on other Electronic Products offered by the Society, contact:

Electronic Products and Services Phone: 800-321-4267 or 401-455-4166

Internet: eps@math.ams.org

To obtain AMS-TEX, AMS-L^{A}TEX, or AMSFonts on diskette contact our Customer Services Department. There is a fee of approximately $15–$25 for each item ordered. You will need to specify whether you want Macintosh or IBM compatible diskettes and in the case of IBM compatible systems both the size and density of the diskettes.

Customer Services Phone: 800-321-4267

Internet: cust-serv@math.ams.org

Author Packages

Code  Full Series Name
bull  Bulletin (New Series) of the American Mathematical Society
cbmath CBMS Issues in Mathematics Education
cbms  CBMS Regional Conference Series in Mathematics
cmsams Conference Proceedings, Canadian Mathematical Society
coll  Colloquium Publications
conm-m Contemporary Mathematics (Monograph)
conm-p Contemporary Mathematics (Proceedings)
crmm  CRM Monograph Series
crmp  CRM Proceedings & Lecture Notes
dimacs DIMACS: Series in Discrete Mathematics and Theoretical Computer Science
fic  Fields Institute Communications
fim  Fields Institute Monographs
gsm  Graduate Studies in Mathematics
hmath History of Mathematics
jams  Journal of the American Mathematical Society
lam  Lectures in Applied Mathematics
llsci Lectures on Mathematics in the Life Sciences
mcom  Mathematics of Computation
memo  Memoirs of the American Mathematical Society
mmono Translations of Mathematical Monographs
pcms  IAS/Park City Mathematics Series
proc  Proceedings of the American Mathematical Society
psapm Proceedings of Symposia in Applied Mathematics
pspum Proceedings of Symposia in Pure Mathematics
suga  Sugaku Expositions
surv  Mathematical Surveys and Monographs
tran  Transactions of the American Mathematical Society
ulect University Lecture Series

The most up-to-date list of available author packages is in the READ.ME file in the FTP e-MATH directory /pub/author-info.

The AMS has created style files which can be used if you are preparing a paper or monograph and are uncertain to which publication you will be submitting. Journals: use

• xxx-j.sty for AMS-TEX;

• xxx-j-l.cls for AMS-L^{A}TEX Book proceedings: use

• xxx-p.sty for AMS-TEX;

• xxx-p-l.cls for AMS-L^{A}TEX Book monographs: use

• xxx-m.sty for AMS-TEX;

• xxx-m-l.cls for AMS-L^{A}TEX